# Providence, USA

### v1

@@ Abstract
We investigate the phenomenology of a dark QCD sector interacting with the Standard Model (SM) via the electroweak (EW) portals. The portal interactions allow SM bosons, such as $Z$ and $h$, or additional bosons that mix with them, to decay into dark quarks, producing dark showers. The light dark mesons are expected to be long-lived particles (LLPs), as their decays back to the SM states through the EW-portal interactions typically have macroscopic decay lengths. We focus on dark shower events initiated by various bosons at the Large Hadron Collider (LHC). The most prominent signal is the displaced decay of GeV-scale dark pions as LLPs. Current limits on dark shower signals at LHC detectors are recast from public data in a model-independent manner. Future limits in the high-luminosity phase and proposed auxiliary detectors are also projected. Additionally, we study the flavor-changing neutral current (FCNC) $B$ decays into dark pions, obtaining both current and projected constraints at the LHC and other facilities. These constraints can be combined for specific models, which are illustrated in two EW-portal benchmarks: one with the heavy doublet fermion mediation and another with the $Z^{\prime}$ mediator including a mass mixing. The collider reach shows significant potential to probe the parameter space unconstrained by EW precision tests, highlighting the necessity of dedicated LLP search strategies and facilities.
@@ 1 Introduction
A dark sector containing a quantum chromodynamics (QCD)-like structure which interacts very weakly with the Standard Model (SM) appears in many new physics scenarios, motivated by important questions of nature, e.g., the neutral naturalness Chacko:2005pe ; Burdman:2006tz ; Cai:2008au ; Cheng:2018gvu ; Cohen:2018mgv or cosmological relaxation Graham:2015cka solutions to the hierarchy problem. It also provides possible candidates of the dark matter (DM) with dark hadrons if they are cosmologically stable. Experimentally, if the dark quarks or gluons are produced at energies much higher than the confining scale, such as in a collider environment, they are expected to generate dark showers and end up with multiple dark hadrons. Depending on whether these dark hadrons decay back to SM particles and their lifetimes, they can give rise to a variety of exotic collider signals, including displaced decays of long-lived particles (LLPs) Strassler:2006im ; Strassler:2006qa ; Han:2007ae ; Alimena:2019zri , emerging jets Schwaller:2015gea , semi-visible jets Cohen:2015toa , etc. Therefore, these models are also of great experimental interest to explore.
@@
The collider phenomenology of the dark QCD depends on the portal interactions between the SM and the dark sectors, which govern both dark particle production and dark hadron decays. If the dark sector and the SM only couple through some very heavy ($\gtrsim$ TeV) states, at low energies the mediators can be integrated out to induce higher-dimensional effective interactions between singlet SM operators and dark operators. Without additional light mediators, it is expected that a significant fraction (if not all) of the light dark hadrons is long-lived or stable. If the SM singlet operator that couples to the dark sector always involves two or more SM particles, the production of the dark particles is enhanced at energies near the mediator scale. At the Large Hadron Collider (LHC), the dark particles are then typically produced with high energies, including the possibilities of associated energetic visible particles or jets, which could help trigger the events. The dark showers induced by the dark quarks or gluons can give rise to a variety of experimental signatures, depending on the dark QCD dynamics and the portal interactions. They produce semi-visible jets if only a fraction of the dark hadrons give visible decays. If the dark hadrons have displaced decays within the detector, the corresponding jets will appear as emerging jets. Searches for prompt semi-visible jets have put strong bounds on the scales that suppress higher dimensional operators or the heavy mediator masses Cohen:2017pzm ; CMS:2021dzg ; ATLAS:2023swa . For the $s$-channel $Z^{\prime}$ mediator with a coupling 0.25 to SM quarks, the $Z^{\prime}$ mass is excluded in the range 1.5–5.1 TeV for the invisible fraction $0.780.01<r_{\rm inv}<$ 0.78 CMS:2021dzg . For the $t$-channel model where the mediator couples to an SM quark and a dark quark, the mediator mass is constrained to be above 2.4–2.7 TeV for a coupling $\lambda=1$ and $0.90.1<r_{\rm inv}<$ 0.9 ATLAS:2023swa . The existing emerging jet search is based on a specific model where a heavy mediator particle decays to an SM jet and an emerging jet CMS:2018bvr ; CMS:2024rea . The heavy mediator mass is excluded up to 1.95 TeV for the dark hadron decay length less than 100 mm, and 1.2 TeV for decay length $\sim 1$ m. If all dark hadrons have long lifetimes and are stable at the collider scale, the dark showers will be invisible, and the collider bounds come from mono-X searches just as the usual DM searches. For an $s$-channel vector mediator, the monojet searches have put a $\sim 2$ TeV bound on the $Z^{\prime}$ mass with a coupling 0.25 to SM quarks CMS:2021far ; ATLAS:2021kxv .
@@
The kinematics of the dark showers initiated by $Z$, $h$ (and $Z^{\prime}$, $\phi$) depends only on the parameters of the dark sector but not the mediators in contrast to high-scale mediator models. The event rates can be effectively parametrized by the exotic branching ratios of $Z$ and $h$, or the mixing parameters of $Z^{\prime}$ and $\phi$. Therefore, dark shower searches in these cases can be performed in a model-independent way, with only a few additional parameters like the dark hadron masses, their decay lifetimes and modes. Such model-independent results allow easy reinterpretations of constraints or reaches of specific models. A major motivation of this work is to provide such model-independent evaluations of dark shower searches from EW-scale $Z,Z^{\prime},h,\phi$ decays at various detectors, including current constraints and future projections.
@@
If the dark hadrons are light enough to appear in the flavor-changing neutral current (FCNC) decays of SM hadrons, the FCNC decays can also provide probes of the dark sector Kamenik:2011vy . Even if the portal interactions conserve SM flavor, the FCNC amplitudes still inevitably arise at one loop. As the widths of SM FCNC decays are small, the FCNC processes are sensitive to other tiny exotic decays beyond the SM. Furthermore, heavy-flavored QCD hadrons have masses significantly below the EW scale, which implies that not only high-energy experiments but also high-intensity experiments can produce a substantial number of heavy-flavored hadrons Kou:2018nap . The exotic decays of the heavy-flavored hadrons are studied in various experiments CMS:2016plw ; LHCb:2020frr ; Belle-II:2023esi ; CHARM:1985anb . They also allow a model-independent interpretation of the branching ratios to the dark sector, and the lifetimes and decay modes of the dark hadrons. They have different dependences on the dark sector parameters relative to the dark showers from $Z,Z^{\prime},h,\phi$ decays, so they provide complementary tests. The comparisons of the experimental reaches from $Z,Z^{\prime},h$, and $\phi$ decays and FCNC processes can be made within specific benchmark models.
@@
The rest of the paper is organized as follows. In Sec. 2, we briefly introduce the phenomenology of dark hadrons with EW-portal couplings, including their composition, symmetry, and the effective field theory (EFT) of decays. Sec. 3 introduces various collider constraints from dark shower production with LLP signals. The results are shown in a model-independent way, including both recasts of current data and projections for future experiments. Similarly, analyses and constraints for dark hadron production in FCNC processes are presented in Sec. 4 in a model-independent way. The above model-independent results are combined according to a couple benchmark models in Sec. 5, where different production mechanisms are related by UV parameters. We briefly summarize in Sec. 6. App. A describes the dimuon vertex trigger efficiency that we extrapolate for HL-LHC projections. App. B contains the details of the two benchmark models used to compare the reaches of various experiments.
@@ 2 Phenomenology of Dark Hadrons
Above the dark pions, there are vector mesons ($\hat{\rho}$) in the dark hadron spectrum. They can be interesting if they are not much heavier than the dark pions and their decays to dark pions are forbidden. In that case, they may also decay back to SM states, constituting part of the dark shower. The decay of dark vector mesons in the EW-portal case happens through their mixing with the SM photon or $Z$ boson, with the latter case a characteristic phenomenon of the EW portal. The branching ratios to SM states for a dark vector meson decay through the $Z$ mixing are shown in Fig. 2, based on the method of Ref. Baruch:2022esd . Since when $m_{\hat{\rho}}>2m_{\hat{\pi}}$, all $\hat{\rho}$ decay to $\hat{\pi}$ and leave no significant impact, for the model-independent limits we focus on the dark showers made of dark pions. The results can be re-interpreted for the dark vector mesons if one has the knowledge of the multiplicity of the dark vector mesons in the dark shower.
@@
3 Model-Independent Limits on Dark Shower Signals
With the EW portal, we consider dark showers at a high energy collider generated from the decays of particles at the EW scale, including SM $Z$, $h$, and also possible dark $Z^{\prime}$ and scalar $\phi$ which mostly couple to the dark sector. At $\sqrt{s}=13(14)$ TeV, the total production rates of $Z$ and $h$ from $pp$ collisions are LHCHiggsCrossSectionWorkingGroup:2016ypw ; Ogul:2017zjd ; ATLAS:2024nrd
@@ 3 Model-Independent Limits on Dark Shower Signals
Figure 3: Cross section of $Z^{\prime}$ and $\phi$ production at 13.6 TeV $pp$ collisions as a function of their masses. LEFT: $Z^{\prime}$ production with the mixing between $Z$ and $Z^{\prime}$ taken to be $\xi=0.01$. The blue horizontal line represents the cross section of dark quarks from $Z$ production assuming $\text{Br}(Z\to\bar{\psi}\psi)=10^{-4}$. RIGHT: $\phi$ production with the $h$-$\phi$ mixing angle $\theta_{s}=0.1$. For comparison, the blue horizontal line is the dark shower rate from exotic Higgs decays with a BR of 0.01.
@@
The Hidden Valley module Carloni:2010tw ; Carloni:2011kk of Pythia8 Sjostrand:2014zea is employed to simulate various dark shower benchmark samples from heavy boson decays. The dark pion kinetic distributions such as $p_{T}$, $\eta$, and multiplicity of dark hadrons are crucial for our analysis Albouy:2022cin . Fig. 4 displays the average dark pion multiplicities in the dark showers from $Z,Z^{\prime},h,\phi$ decays with different $m_{\hat{\pi}}$. For $m_{\phi}<m_{h}/2$, the light $\phi$ can be produced either from $pp$ collision directly or indirectly from $h\to\phi\phi$ decays. For all benchmark EW-portal models, the average dark pion multiplicity $N_{\hat{\pi}}$ drops with $m_{\hat{\pi}}$ with $N_{\hat{\pi}}\propto m_{\hat{\pi}}^{-0.6}$ approximately. The simulated dark pion kinematic distributions from various initial vector boson benchmarks are shown in Fig. 5, while all scalar boson cases are plotted in Fig 6. In all cases, the majority of dark pions are soft ($p_{T}\lesssim 10$ GeV), while the high $p_{T}$ tail strongly depends on the initial boson mass. Since the initial bosons are also moderately boosted by QCD emission, the maximum dark pion $p_{T}$ may exceed half of the initial boson mass. The small average dark pion $p_{T}$ thus reduces the signal efficiency for searches only sensitive to hard tracks. The average $|\eta|$ of dark pions decreases with increasing the initial boson mass. Also, dark pions produced from $Z^{(\prime)}$ have higher $|\eta|$ when compared to those from scalar bosons with a similar mass. This is because the major production channel for $Z^{(\prime)}$ is $q\bar{q}$ with highly asymmetric parton distribution functions (PDFs), in contrast to the dominant $gg$ fusion in the scalar cases. In practice, the $\eta$ distribution in the high $|\eta|$ tail is directly related to the performance of experiments like LHCb. We also plot the spread of dark pion $p_{T}$ within a single event in Fig. 5. In particular, the plotted dimensionless quantity is the standard deviation of dark pion $p_{T}$ divided by their average in each event. Generically, as the initial boson mass increases, generating more dark pions in the event, the dark pion $p_{T}$ are more unevenly distributed in the event.
@@ 3.1 Dark Shower Sensitivity at CMS
Given that the analysis of Ref. CMS:2021sch targets different models, it is essential to recast the publicly available data for dark shower searches. The procedure follows a similar approach in Ref. Born:2023vll . A set of truth-level cuts is first applied on each dimuon pair resulting from dark hadron decays. Only the muon pairs with both $p_{T,\mu}>3$ GeV, $|\eta_{\mu}|<2.4$ are selected. In addition, the DV’s transverse displacement $l_{xy}$ must be $<$11 cm. The angle ($\Delta\phi_{\hat{\pi}}$) between the DV’s momentum and its spatial displacement from the primary vertex (PV) must be less than 0.02. When $m_{\hat{\pi}}$ is small, the DV’s invariant mass window primarily results from the finite resolution of the tracker, which is about 1.1% of the resonance mass. We therefore take the window to be $m_{\mu\mu}\in[0.978m_{\hat{\pi}},1.022m_{\hat{\pi}}]$ to maximize sensitivity. [^2] Due to final state photon radiation, the reconstructed $m_{\mu\mu}$ sometimes lie outside the window, further reducing the signal efficiency. The effect is more significant for heavier $m_{\hat{\pi}}$, leading to a maximum signal efficiency loss $\sim 25\%$, which is included in the final results. To obtain the final signal yield, each remaining candidate DV will be weighed by the trigger and selection efficiencies reported in Ref. CMS:2021sch according to their properties, such as their $p_{T}$ or $l_{xy}$. As dark pions are likely to decay back to SM particles within a small solid angle, no isolation requirements are placed on candidate DVs. [^3] A similar analysis in Ref. Born:2023vll demonstrates that the isolation requirements on muons will not significantly affect the final exclusion limit. The signal yield is divided into several $l_{xy}$ bins. Typically, the results are dominated by one bin so only the maximal signal significance is adopted to get the model-independent limit.
For a given $m_{\hat{\pi}}$, the background is calculated based on the average background level and the width of the $m_{\mu\mu}$ window since backgrounds follow an approximate uniform-distribution around each narrow window. To better match with the reported maximum signal yields reported in Ref. CMS:2021sch , the background yield in each $l_{xy}$ bin is assumed to have a systematic uncertainty. A relative systematic uncertainty of 3% in each bin leads to results compatible with the signal event upper limit reported by CMS.
@@
Figure 7: Model-independent limits on $Z$, $Z^{\prime}$, $h$, and $\phi$-initiated dark shower by recasting current CMS dimuon DV search CMS:2021sch . The different colors correspond to several $m_{\hat{\pi}}$ benchmark values. All limits are modified by BR$(\hat{\pi}\to\mu^{+}\mu^{-})$. TOP-LEFT: The current CMS limits on the product of the exotic BR$(Z\to\psi\bar{\psi})$. TOP-RIGHT: The current limit of a $Z^{\prime}$ that mainly decays to dark shower with $m_{Z^{\prime}}=20$ GeV. The limits are given in terms of $\xi^{2}$. BOTTOM-LEFT: The limit on exotic Higgs BR($h\to\psi\bar{\psi}$). BOTTOM-RIGHT: The current limit of the light scalar $\phi$ that mainly decays to dark shower with $m_{\phi}=70$ GeV. The limits are given in terms of its Higgs mixing angle squared $\theta_{s}^{2}$.
All limits are delivered model independently as functions of $c\tau(\hat{\pi})$ and $m_{\hat{\pi}}$ only. For the SM initial bosons $Z$ and $h$, the limits are set on their exotic BR to dark shower. Limits from BSM initial bosons $Z^{\prime}$ and $\phi$ are given as their mixing parameters $\xi^{2}$ or $\theta_{s}^{2}$. The limits from all four cases are scaled by BR($\hat{\pi}\to\mu^{+}\mu^{-}$) due to the nature of 1DV searches. Conversely, if more than one dimuon DVs with the same mass are found in a single event, backgrounds from flavored hadrons and detector effects shall diminish as it is unlikely to give two DVs with the same mass this way. However, with the data scouting approach, the muon $p_{T}>3$ GeV cut leads to low DV efficiency of $\mathcal{O}(10^{-3})$. The chances of reconstructing more than one DV in an event are too low to give useful constraints with the current data, but could be relevant at the HL-LHC discussed later.

### v2

@@ Abstract
We investigate the phenomenology of a dark QCD sector interacting with the Standard Model (SM) via the electroweak (EW) portals. The portal interactions allow SM bosons, such as $Z$ and $h$, or additional bosons that mix with them, to decay into dark quarks, producing dark showers. The light dark mesons are expected to be long-lived particles (LLPs), as their decays back to the SM states through the EW-portal interactions typically have macroscopic decay lengths. We focus on dark shower events initiated by various bosons at the Large Hadron Collider (LHC). The most prominent signal is the displaced decay of GeV-scale dark pions as LLPs. Current limits on dark shower signals at LHC detectors are recast from public data to provide simplified limits insensitive to UV physics details. Future limits in the high-luminosity phase and proposed auxiliary detectors are also projected. Additionally, we study the flavor-changing neutral current (FCNC) $B$ decays into dark pions, obtaining both current and projected constraints at the LHC and other facilities. These constraints can be combined for specific models, which are illustrated in two EW-portal benchmarks: one with the heavy doublet fermion mediation and another with the $Z^{\prime}$ mediator including a mass mixing. The collider reach shows significant potential to probe the parameter space unconstrained by EW precision tests, highlighting the necessity of dedicated LLP search strategies and facilities.
@@ 1 Introduction
A dark sector containing a quantum chromodynamics (QCD)-like structure which interacts very weakly with the Standard Model (SM) appears in many new physics scenarios, motivated by important questions of nature, e.g., the neutral naturalness Chacko:2005pe ; Burdman:2006tz ; Cai:2008au ; Cheng:2018gvu ; Cohen:2018mgv or cosmological relaxation Graham:2015cka solutions to the hierarchy problem. It also provides possible candidates of the dark matter (DM) with dark hadrons if they are cosmologically stable. Experimentally, if the dark quarks or gluons are produced at energies much higher than the confining scale, such as in a collider environment, they are expected to generate dark showers and end up with multiple dark hadrons. Depending on whether these dark hadrons decay back to SM particles and their lifetimes, they can give rise to a variety of exotic collider signals, including displaced decays of long-lived particles (LLPs) Strassler:2006im ; Strassler:2006qa ; Han:2007ae ; Alimena:2019zri , emerging jets Schwaller:2015gea , semi-visible jets Cohen:2015toa , etc. Therefore, these models are also of great experimental interest to explore  [^1] See also Ref. Park:2017rfb for related phenomenology..
@@
The collider phenomenology of the dark QCD depends on the portal interactions between the SM and the dark sectors, which govern both dark particle production and dark hadron decays. If the dark sector and the SM only couple through some very heavy ($\gtrsim$ TeV) states, at low energies the mediators can be integrated out to induce higher-dimensional effective interactions between singlet SM operators and dark operators. Without additional light mediators, it is expected that a significant fraction (if not all) of the light dark hadrons is long-lived or stable. If the SM singlet operator that couples to the dark sector always involves two or more SM particles, the production of the dark particles is enhanced at energies near the mediator scale. At the Large Hadron Collider (LHC), the dark particles are then typically produced with high energies, including the possibilities of associated energetic visible particles or jets, which could help trigger the events. The dark showers induced by the dark quarks or gluons can give rise to a variety of experimental signatures, depending on the dark QCD dynamics and the portal interactions. They produce semi-visible jets if only a fraction of the dark hadrons give visible decays. If the dark hadrons have displaced decays within the detector, the corresponding jets will appear as emerging jets. Searches for prompt semi-visible jets have put strong bounds on the scales that suppress higher dimensional operators or the heavy mediator masses Cohen:2017pzm ; CMS:2021dzg ; ATLAS:2023swa . For the $s$-channel $Z^{\prime}$ mediator with a coupling 0.25 to SM quarks, the $Z^{\prime}$ mass can be constrained for the invisible fraction $0.780.01<r_{\rm inv}<$ 0.78. It is excluded from 1.5 TeV up to 5.1 TeV for $r_{\rm inv}$ around 0.3 and between 2.1 TeV and 3.3 TeV for $r_{\rm inv}\sim 0.7$ CMS:2021dzg . For the $t$-channel model where the mediator couples to an SM quark and a dark quark, the mediator mass is constrained to be above 2.4–2.7 TeV for a coupling $\lambda=1$ and $0.90.1<r_{\rm inv}<$ 0.9 ATLAS:2023swa . The existing emerging jet search is based on a specific model where a heavy mediator particle decays to an SM jet and an emerging jet CMS:2018bvr ; CMS:2024rea . The heavy mediator mass is excluded up to 1.95 TeV for the dark hadron decay length less than 100 mm, and 1.2 TeV for decay length $\sim 1$ m. If all dark hadrons have long lifetimes and are stable at the collider scale, the dark showers will be invisible, and the collider bounds come from mono-X searches just as the usual DM searches. For an $s$-channel vector mediator, the monojet searches have put a $\sim 2$ TeV bound on the $Z^{\prime}$ mass with a coupling 0.25 to SM quarks CMS:2021far ; ATLAS:2021kxv .
@@
The kinematics of the dark showers initiated by $Z$, $h$ (and $Z^{\prime}$, $\phi$) depends only on the parameters of the dark sector but not the mediators in contrast to high-scale mediator models. The event rates can be effectively parametrized by the exotic branching ratios of $Z$ and $h$, or the mixing parameters of $Z^{\prime}$ and $\phi$. Therefore, dark shower searches in these cases can be performed in a simplified approach with only a few additional parameters like the dark hadron masses, their decay lifetimes and modes. Such results allow easy reinterpretations of constraints or reaches of specific UV models. A major motivation of this work is to provide rough evaluations of dark shower searches from EW-scale $Z,Z^{\prime},h,\phi$ decays at various detectors, including current constraints and future projections. Note that the limits obtained this way will always succumb to large systematics from unknown dark shower dynamics (see, $e.g.$, Ref. Cohen:2020afv ; Albouy:2022cin ). The fact doesn’t contradict our goal of providing quick projections.
@@
If the dark hadrons are light enough to appear in the flavor-changing neutral current (FCNC) decays of SM hadrons, the FCNC decays can also provide probes of the dark sector Kamenik:2011vy . Even if the portal interactions conserve SM flavor, the FCNC amplitudes still inevitably arise at one loop. As the widths of SM FCNC decays are small, the FCNC processes are sensitive to other tiny exotic decays beyond the SM. Furthermore, heavy-flavored QCD hadrons have masses significantly below the EW scale, which implies that not only high-energy experiments but also high-intensity experiments can produce a substantial number of heavy-flavored hadrons Kou:2018nap . The exotic decays of the heavy-flavored hadrons are studied in various experiments CMS:2016plw ; LHCb:2020frr ; Belle-II:2023esi ; CHARM:1985anb . This also allows us to perform simplified projections of dark sector reach. The relevant parameters will be the FCNC branching ratios to the dark sector, and the lifetimes and decay modes of the dark hadrons. These projections have different dependences on the dark sector parameters relative to the dark showers from $Z,Z^{\prime},h,\phi$ decays, so they provide complementary tests. The comparisons of the experimental reaches from $Z,Z^{\prime},h$, and $\phi$ decays and FCNC processes can be made within specific benchmark models.
@@
The rest of the paper is organized as follows. In Sec. 2, we briefly introduce the phenomenology of dark hadrons with EW-portal couplings, including their composition, symmetry, and the effective field theory (EFT) of decays. Sec. 3 introduces various collider constraints from dark shower production with LLP signals. The results are shown, including both recasts of current data and projections for future experiments. Similarly, analyses and constraints for dark hadron production in FCNC processes are presented in Sec. 4. The above results are combined according to a couple benchmark models in Sec. 5, where different production mechanisms are related by UV parameters. We briefly summarize in Sec. 6. App. A describes the dimuon vertex trigger efficiency that we extrapolate for HL-LHC projections. App. B contains the details of the two benchmark models used to compare the reaches of various experiments.
@@ 2 Phenomenology of Dark Hadrons
Above the dark pions, there are vector mesons ($\hat{\rho}$) in the dark hadron spectrum. They can be interesting if they are not much heavier than the dark pions and their decays to dark pions are forbidden. In that case, they may also decay back to SM states, constituting part of the dark shower. The decay of dark vector mesons in the EW-portal case happens through their mixing with the SM photon or $Z$ boson, with the latter case a characteristic phenomenon of the EW portal. The branching ratios to SM states for a dark vector meson decay through the $Z$ mixing are shown in Fig. 2, based on the method of Ref. Baruch:2022esd . Since when $m_{\hat{\rho}}>2m_{\hat{\pi}}$, all $\hat{\rho}$ decay to $\hat{\pi}$ and leave no significant impact, for these limits we focus on the dark showers made of dark pions. The results can be re-interpreted for the dark vector mesons if one has the knowledge of the multiplicity of the dark vector mesons in the dark shower.
@@
3 Limits on Dark Shower Signals
Throughout the paper, we assume a simple dark pion benchmark with two light dark flavors and thus three dark pions. An $SU(3)$ confining dark gauge interaction is assumed, with a gauge coupling of 0.1 around the $Z$ mass. With the EW portal, we consider dark showers at a high energy collider generated from the decays of particles at the EW scale, including SM $Z$, $h$, and also possible dark $Z^{\prime}$ and scalar $\phi$ which mostly couple to the dark sector. In this work, we set the simulation for (HL-)LHC with $\sqrt{s}=13(14)$ TeV, respectively. The total production rates of $Z$ and $h$ from $pp$ collisions are LHCHiggsCrossSectionWorkingGroup:2016ypw ; Ogul:2017zjd ; ATLAS:2024nrd
@@ 3 Limits on Dark Shower Signals
Figure 3: Cross section of $Z^{\prime}$ and $\phi$ production at 14 TeV $pp$ collisions as a function of their masses. LEFT: $Z^{\prime}$ production with the mixing between $Z$ and $Z^{\prime}$ taken to be $\xi=0.01$. The blue horizontal line represents the cross section of dark quarks from $Z$ production assuming $\text{Br}(Z\to\bar{\psi}\psi)=10^{-4}$. RIGHT: $\phi$ production with the $h$-$\phi$ mixing angle $\theta_{s}=0.1$. For comparison, the blue horizontal line is the dark shower rate from exotic Higgs decays with a BR of 0.01.
@@
The Hidden Valley module Carloni:2010tw ; Carloni:2011kk of Pythia8 Sjostrand:2014zea is employed to simulate various dark shower benchmark samples from heavy boson decays. The dark hadronization is simulated with minimal dark pion multiplicity. [^2] The choice corresponds to changing the dimensionless parameters aLund to 0.1 and bmqv2 to 2.0 in the Pythia8 Hidden Valley module. The other parameters are kept as their default values. Though the smaller dark pion multiplicity from this benchmark value reduces the overall signal rate, the overall signal efficiency is compensated by the fact that the average dark pion energy becomes higher, leaving the final result less sensitive to these parameters. The dark pion kinetic distributions such as $p_{T}$, $\eta$, and multiplicity of dark hadrons are crucial for our analysis Albouy:2022cin . Fig. 4 displays the average dark pion multiplicities in the dark showers from $Z,Z^{\prime},h,\phi$ decays with different $m_{\hat{\pi}}$. For $m_{\phi}<m_{h}/2$, the light $\phi$ can be produced either from $pp$ collision directly or indirectly from $h\to\phi\phi$ decays. For all benchmark EW-portal models, the average dark pion multiplicity $N_{\hat{\pi}}$ drops with $m_{\hat{\pi}}$ with $N_{\hat{\pi}}\propto m_{\hat{\pi}}^{-0.6}$ approximately. The simulated dark pion kinematic distributions from various initial vector boson benchmarks are shown in Fig. 5, while all scalar boson cases are plotted in Fig 6. In all cases, the majority of dark pions are soft ($p_{T}\lesssim 10$ GeV), while the high $p_{T}$ tail strongly depends on the initial boson mass. Since the initial bosons are also moderately boosted by QCD emission, the maximum dark pion $p_{T}$ may exceed half of the initial boson mass. The small average dark pion $p_{T}$ thus reduces the signal efficiency for searches only sensitive to hard tracks. The average $|\eta|$ of dark pions decreases with increasing the initial boson mass. Also, dark pions produced from $Z^{(\prime)}$ have higher $|\eta|$ when compared to those from scalar bosons with a similar mass. This is because the major production channel for $Z^{(\prime)}$ is $q\bar{q}$ with highly asymmetric parton distribution functions (PDFs), in contrast to the dominant $gg$ fusion in the scalar cases. In practice, the $\eta$ distribution in the high $|\eta|$ tail is directly related to the performance of experiments like LHCb. We also plot the spread of dark pion $p_{T}$ within a single event in Fig. 5. In particular, the plotted dimensionless quantity is the standard deviation of dark pion $p_{T}$ divided by their average in each event. Generically, as the initial boson mass increases, generating more dark pions in the event, the dark pion $p_{T}$ are more unevenly distributed in the event.
@@ 3.1 Dark Shower Sensitivity at CMS
Given that the analysis of Ref. CMS:2021sch targets different models, it is essential to recast the publicly available data for dark shower searches. The procedure follows a similar approach in Ref. Born:2023vll . A set of truth-level cuts is first applied on each dimuon pair resulting from dark hadron decays. Only the muon pairs with both $p_{T,\mu}>3$ GeV, $|\eta_{\mu}|<2.4$ are selected. In addition, the DV’s transverse displacement $l_{xy}$ must be $<$11 cm. The angle ($\Delta\phi_{\hat{\pi}}$) between the DV’s momentum and its spatial displacement from the primary vertex (PV) must be less than 0.02. When $m_{\hat{\pi}}$ is small, the DV’s invariant mass window primarily results from the finite resolution of the tracker, which is about 1.1% of the resonance mass. We therefore take the window to be $m_{\mu\mu}\in[0.978m_{\hat{\pi}},1.022m_{\hat{\pi}}]$ to maximize sensitivity. [^4] Due to final state photon radiation, the reconstructed $m_{\mu\mu}$ sometimes lie outside the window, further reducing the signal efficiency. The effect is more significant for heavier $m_{\hat{\pi}}$, leading to a maximum signal efficiency loss $\sim 25\%$, which is included in the final results. To obtain the final signal yield, each remaining candidate DV will be weighed by the trigger and selection efficiencies reported in Ref. CMS:2021sch according to their properties, such as their $p_{T}$ or $l_{xy}$. As dark pions are likely to decay back to SM particles within a small solid angle, no isolation requirements are placed on candidate DVs. [^5] A similar analysis in Ref. Born:2023vll demonstrates that the isolation requirements on muons will not significantly affect the final exclusion limit. The signal yield is divided into several $l_{xy}$ bins. Typically, the results are dominated by one bin, so only the maximal signal significance from the dominating $l_{xy}$ bin is adopted to get the limit instead of combining different bins.
Since the CMS reported backgrounds change very slowly with $m_{\mu\mu}$ away from SM resonances, we calculate the background yield based on the average background yield per $m_{\mu\mu}$ provided and the $m_{\mu\mu}$ mass window defined above. To better match with the reported maximum signal yields reported, the background yield in each $l_{xy}$ bin is assumed to have a systematic uncertainty. A 3% relative background fluctuation in each $l_{xy}$ bin is assumed, which leads to results compatible with the signal event upper limit reported by CMS. The 95% C.L. limits on the exotic BRs are obtained by solving the approximate relation
$\frac{S_{95}}{\sqrt{S_{95}+B+(\Delta B)^{2}}}=2~{},$ (8)
where $S_{95}$, $B$ and $\Delta B$ stand for the signal yield at 95% C.L., the background yield and the systematic background uncertainty in the dominating $l_{xy}$ bin, respectively.
@@
Figure 7: The 95% C.L. limits on $Z$, $Z^{\prime}$, $h$, and $\phi$-initiated dark shower by recasting current CMS dimuon DV search CMS:2021sch . The different colors correspond to several $m_{\hat{\pi}}$ benchmark values. All limits are modified by BR$(\hat{\pi}\to\mu^{+}\mu^{-})$. The kinks in the small $c\tau$ region are induced by the dominant $l_{xy}$ bin switch between 0.2-1 and 1-2.4 cm, which are of different behaviors. TOP-LEFT: The current CMS limits on the product of the exotic BR$(Z\to\psi\bar{\psi})$. TOP-RIGHT: The current limit of a $Z^{\prime}$ that mainly decays to dark shower with $m_{Z^{\prime}}=20$ GeV. The limits are given in terms of $\xi^{2}$. BOTTOM-LEFT: The limit on exotic Higgs BR($h\to\psi\bar{\psi}$). BOTTOM-RIGHT: The current limit of the light scalar $\phi$ that mainly decays to dark shower with $m_{\phi}=70$ GeV. The limits are given in terms of its Higgs mixing angle squared $\theta_{s}^{2}$.
All limits are delivered as functions of $c\tau(\hat{\pi})$ and $m_{\hat{\pi}}$ only. For the SM initial bosons $Z$ and $h$, the limits are set on their exotic BR to dark shower. Limits from BSM initial bosons $Z^{\prime}$ and $\phi$ are given as their mixing parameters $\xi^{2}$ or $\theta_{s}^{2}$. The limits from all four cases are scaled by BR($\hat{\pi}\to\mu^{+}\mu^{-}$) due to the nature of 1DV searches. Conversely, if more than one dimuon DVs with the same mass are found in a single event, backgrounds from flavored hadrons and detector effects shall diminish as it is unlikely to give two DVs with the same mass this way. However, with the data scouting approach, the muon $p_{T}>3$ GeV cut leads to low DV efficiency of $\mathcal{O}(10^{-3})$. The chances of reconstructing more than one DV in an event are too low to give useful constraints with the current data, but could be relevant at the HL-LHC discussed later.